# Apples to Apples: Establishing Comparability in Knowledge Generation Tasks Involving Users

### v1

@@ 4 Results and Analysis
4.0.1 Reliability and Internal Consistency
Reliability refers to the degree to which the items within a test or survey consistently measure the same construct. High internal consistency strengthens the statistical reliability of metrics, thereby enhancing the validity of group comparisons. To evaluate internal consistency, we recommend using Cronbach’s Alpha. Higher alpha coefficients indicate greater shared covariance among items, suggesting they assess the same underlying concept. A Cronbach’s Alpha value of $\geq 0.7$ is generally considered acceptable. [24]
4.0.2 Data Normality
Data normality refers to how much data distribution aligns with a normal curve. While normality is not always required, t-tests and ANOVA that assume data follows such a distribution. ANOVA is fairly robust when data is not normality distributed when sample sizes are large, but that is difficult when dealing with user studies. We, therefore, require studies to test for normality and report on normality. Participants may, if they wish, avail of other statistical measures. As the sample sizes of a group will likely not exceed 50, we propose the Shapiro-Wilk test to assess normality. In this test, the sample is compared to a theoretical normal distribution.
4.0.3 Homogeneity
@@
4.0.4 Group Comparisons
To determine whether differences between groups are statistically significant, researchers should employ well-established statistical tests. The choice of test depends on the assumptions about the data. The recommended tests when the data is normal (also known as parametric tests) are Welch’s t-test when comparing two groups and ANOVA when comparing more than two groups simultaneously. Both tests assume normality and homogeneity of variance. The recommended tests when the data is not normal (also known as non-parametric tests) are the Wilcoxon test for comparing two groups and the Kruskal-Wallis test for multiple.
4.0.5 Correlation Analysis
Correlation methods assess the strength and direction of the relationship between variables, which can provide deeper insights into study outcomes. For instance, examining correlations between usability, accuracy, and mental workload can reveal important relationships in user behavior. When data is normally distributed, we recommend the Pearson’s Correlation to measure the strength of a relationship, for example, between accuracy and usability. Otherwise, one should use the Spearman’s Correlation. Researchers must report on correlations between all relevant variables (e.g., usability and mental workload, usability and accuracy, etc.) to provide a comprehensive analysis.
4.0.6 Transparency and Accessibility
To promote transparency and reproducibility, all collected data and the statistical tests performed should be made publicly available online. Providing access to raw data, analysis scripts, and detailed methodology facilitates validation and enhances the study’s credibility. Moreover, sharing such data would allow one to more easily compare results across studies, provided that the conditions are similar.
@@ 5 Discussion
We presented a comparison of user studies in the KGC domain, and we noticed that all studies were different. This makes it impossible to compare KGC languages, tools, and software. To this end, we analyzed the related work, distilled elements we appreciated, and proposed others to establish a common protocol. As such, we proposed a resource, a user study protocol, that provides the KG generation community with a better way to present, compare, and scrutinize contributions.
When designing the protocol, we selected and refined elements from the state-of-the-art that we appreciated. Examples include the use of accuracy and task execution time as simple measures, the use of PSSUQ over SUS to obtain more fine-grained information on usability, usefulness, and information quality, and measuring the mental workload right after the task. Reporting on the correlations between the perceived usability, task execution, and mental workload could shed interesting insights. [8] We furthermore provided guidelines on what statistical techniques to use when reporting on user studies with this protocol, as most merely reported on averages.
Somewhat new compared to the related work is our informed decision to adopt an accessible domain, focus on RML-core functionality as a basis, and formulate five tasks. In Section 3, we provided a rationale that, when comparing two groups, the last two tasks can be replaced for one group so that extensions or variants within the same language or tool can be compared. As such, the resource is sufficiently general for use in the KGC community, and the approach in designing this protocol may inspire others within the Semantic Web community.
The resources have not only been made available with a DOI on a long-term preservation platform, but they are also available on a GitHub repository. The latter allows peers to contribute to the project. The directory structure we use allows for variants of the protocol to be made available. [^3] For the reviewers: at the time of writing, we are currently discussing with the W3C KGC Working Group whether the initiative could be moved to https://github.com/kg-construct/.
The protocol has not yet been used at this stage, but its first uses are planned in the spring of 2025. However, many of the protocol’s separate elements are drawn from existing studies. As such, those parts have already been validated in the community. We also aim to engage with the wider KGC community via the W3C working group on adopting this protocol across different institutions.
@@ 6 Conclusions
Prior KGC user studies used different protocols, making comparison impossible. This paper thus highlights the lack of standardized protocols in user studies related to KGC, making it difficult to compare different studies. We present a new protocol to address these inconsistencies, focusing on participant selection, task design, and evaluation metrics. The protocol suggests detailed guidelines for recruiting participants and disclosing potential biases. The process guidelines for informed consent, pre-questionnaires, familiarization activities, task execution, and post-questionnaires. Five specific mapping sub-tasks are proposed, which can be solved with the equivalent of the RML-Core specification. The protocol recommends using the Post-Study System Usability Questionnaire (PSSUQ) for usability, and the NASA Task Load Index (NASA-TLX) for mental workload, among others. We designed the protocol in such a way that one can analyze one group, or compare groups. To this end, we provided guidelines on which statistical instruments to use.
This protocol aims to provide a more comparable evaluation of KGC user studies, ultimately leading to more effective tools for knowledge graph construction. As such, the protocol is an essential artifact for future longitudinal and comparative studies. Future work involved encouraging the adoption of this protocol by various KGC scholars. While ambitious, it is hoped that this protocol will form the basis of a new, open repository of KGC user studies.
{credits}
[…]
The authors have no competing interests to declare relevant to this article’s content.

### v2

@@ 4 Results and Analysis
Reliability and Internal Consistency
Reliability refers to the degree to which the items within a test or survey consistently measure the same construct. High internal consistency strengthens the statistical reliability of metrics, thereby enhancing the validity of group comparisons. We recommend using Cronbach’s Alpha to evaluate internal consistency. Higher alpha coefficients indicate greater shared covariance among items, suggesting they assess the same underlying concept. A Cronbach’s Alpha value of $\geq 0.7$ is generally considered acceptable. [29]
Data Normality
Data normality refers to how much data distribution aligns with a normal curve. While normality is not always required, t-tests and ANOVA assume that data follows such a distribution. ANOVA is relatively robust when data is not normally distributed, when sample sizes are large, but that is difficult when dealing with user studies. We, therefore, require studies to test for normality and report on normality. Participants may, if they wish, use other statistical measures. As the sample sizes of a group will likely not exceed 50, we propose the Shapiro-Wilk test to assess normality. In this test, the sample is compared to a theoretical normal distribution.
Homogeneity
@@
Group Comparisons
To determine whether differences between groups are statistically significant, researchers should employ well-established statistical tests. The choice of test depends on the assumptions about the data. The recommended tests when the data is normally distributed (also known as parametric tests) are Welch’s t-test when comparing two groups and ANOVA when comparing more than two groups simultaneously. Both tests assume normality and homogeneity of variance. The recommended tests when the data is not normally distributed (also known as non-parametric tests) are the Wilcoxon test for comparing two groups and the Kruskal-Wallis test for multiple.
Correlation Analysis
Correlation methods assess the strength and direction of the relationship between variables, which can provide deeper insights into study outcomes. For instance, examining correlations between usability, accuracy, and mental workload can reveal relationships in user behavior. When data is normally distributed, we recommend the Pearson’s Correlation to measure the strength of a relationship, for example, between accuracy and usability. Otherwise, one should use the Spearman’s Correlation. Researchers must report on correlations between all relevant variables (e.g., usability and mental workload, usability and accuracy, etc.) to provide a comprehensive analysis.
Transparency and Accessibility
To promote transparency and reproducibility, the data and the statistical tests performed should be publicly available online, provided that all personally identifiable information is removed and participant anonymity is preserved in accordance with ethical research standards.
Access to data, analysis scripts, and detailed methodology facilitates validation and enhances the study’s credibility. Moreover, sharing such data would allow one to compare results across studies more easily, provided the conditions are similar.
@@ 5 Discussion
We presented a comparison of user studies in the KGC domain, and we noticed that all studies were different. This makes it impossible to compare KGC languages, tools, and software. To this end, we analyzed the related work, distilled elements we appreciated, and proposed others to establish a common protocol. As such, we proposed a resource, a user study protocol, that provides the KGC community with a better way to present, compare, and scrutinize contributions.
When designing the protocol, we selected and refined elements from the state-of-the-art that we appreciated. Examples include the use of accuracy and task execution time as simple measures, the use of PSSUQ over SUS to obtain more fine-grained information on usability, usefulness, and information quality, and measuring the mental workload right after the task. Reporting on the correlations between the perceived usability, task execution, and mental workload could shed interesting insights. [12] We provided guidelines on what statistical techniques to use when reporting on user studies with this protocol, as most merely reported on averages.
Relatively novel compared to the related work is our informed decision to adopt an accessible domain, focus on RML-core functionality as a basis, and formulate five tasks. In Section 3, we provided a rationale that, when comparing two groups, the last two tasks can be replaced for one group so that extensions or variants within the same language or tool can be compared. As such, the resource is sufficiently general for use in the KGC community, and the approach in designing this protocol may inspire others within the Semantic Web community.
The resources have not only been made available with a DOI [^5] https://zenodo.org/records/14507455 on a long-term preservation platform, but they are also available on a GitHub repository. The latter allows peers to contribute to the project. The directory structure we use allows for variants of the protocol to be made available.
The protocol has not yet been used at this stage, but its first use is planned for the spring of 2025. However, many of its separate elements are drawn from existing studies. As such, those parts have already been validated in the community. We also aim to engage with the wider KGC community via the W3C working group on adopting this protocol across different institutions.
@@ 6 Conclusions
Prior KGC user studies used different protocols, making comparison impossible. This paper thus highlights the lack of standardized protocols in user studies related to KGC, making it difficult to compare different studies. We present a new protocol to address these inconsistencies, focusing on participant selection, task design, and evaluation metrics. The protocol suggests detailed guidelines for recruiting participants and disclosing potential biases. The process guidelines for informed consent, pre-questionnaires, familiarization activities, task execution, and post-questionnaires. Five specific mapping tasks are proposed, which can be solved with the equivalent of the RML-Core specification. The protocol recommends using the Post-Study System Usability Questionnaire (PSSUQ) for usability, and the NASA Task Load Index (NASA-TLX) for mental workload, among others. We designed the protocol in such a way that one can analyze one group, or compare groups. To this end, we provided guidelines on which statistical instruments to use.
This protocol aims to provide a more comparable evaluation of KGC user studies, ultimately leading to more effective tools for knowledge graph construction. As such, the protocol is an essential artifact for future longitudinal and comparative studies. We recognize that the proposal has been constructed in a bottom-up fashion for this community, and future work should look into aligning our proposal with methods for comparing the usability of different (information) systems, such as [30]. Finally, future work also involves encouraging the adoption of this protocol by various KGC scholars. While ambitious, it is hoped that this protocol will form the basis of a new, open repository of KGC user studies.
Acknowledgements.
We thank the reviewers for their many (many) thoughtful comments, which greatly improved the paper. Their feedback was invaluable.
Declaration on Generative AI
During the preparation of this work, the author(s) used Grammarly to improve grammar, check spelling, and reword. After using these tool(s)/service(s), the author(s) reviewed and edited the content as needed and take(s) full responsibility for the publication’s content.